OpenStudy (princessaurora):

*Trigonometric Graphs question* Simplify the expression: cos^2(pi/2)-x)/square root1-sin^2(x). *Please help will give medal for this*

zepdrix (zepdrix):

Simplify this?\[\large\rm \frac{\cos^2\left(\frac{\pi}{2}-x\right)}{\sqrt{1-\sin^2(x)}}\]

OpenStudy (princessaurora):

The mathway calculator won't let me square the cos or the sin

zepdrix (zepdrix):

Ahhh sorry I gotta get food out of the oven a sec >.< I can help in a few minutes, unless one of these fellas gets to it first hehe

OpenStudy (princessaurora):

lol okay save me some though

zepdrix (zepdrix):

XD

zepdrix (zepdrix):

Recall that sine and cosine are `cofunctions`. Example of this concept:\[\large\rm \sin(30^o)=\cos(60^o)\]\[\large\rm \sin(70^o)=\cos(20^o)\]We can generalize this by saying:\[\large\rm \sin(x)=\cos(90^o-x)\]Or so it's more useful for us, let's see this relationship in radians,\[\large\rm \sin(x)=\cos\left(\frac{\pi}{2}-x\right)\]

zepdrix (zepdrix):

Do you see how we can apply this relationship to the numerator?\[\large\rm \frac{\color{orangered}{\cos^2\left(\frac{\pi}{2}-x\right)}}{\sqrt{1-\sin^2(x)}}\]

OpenStudy (princessaurora):

Yeah I see

zepdrix (zepdrix):

So what's the top become? :)

OpenStudy (princessaurora):

sin(x)?

zepdrix (zepdrix):

Well, I was showing you this relationship,\[\large\rm \sin(x)=\cos\left(\frac{\pi}{2}-x\right)\]If we square each side,\[\large\rm \sin^2(x)=\cos^2\left(\frac{\pi}{2}-x\right)\]this right side is what we have in our numerator. So I guess we end up with sin^2(x), ya? :)

zepdrix (zepdrix):

\[\large\rm \frac{\color{orangered}{\sin^2(x)}}{\sqrt{1-\sin^2(x)}}\]So that takes care of the top, cool.

OpenStudy (princessaurora):

Alright

zepdrix (zepdrix):

Recall your Pythagorean Identity which involves sine and cosine:\[\large\rm \sin^2(x)+\cos^2(x)=1\]What happens if we subtract sin^2(x) from each side? What do we get? :d Anything useful?

OpenStudy (princessaurora):

is the bottom tan(x)?

zepdrix (zepdrix):

Hmm no :o

OpenStudy (princessaurora):

My answer choices are weird, they go:A) tan(x), B) cos(x) tan(x), C) cos(x) cot(x), D) sin(x) tan(x).

zepdrix (zepdrix):

We're not quite there yet :) we're working on the denominator. When you subtract sin^2(x) from each side, you get something like this:\[\large\rm \cos^2(x)=\color{orangered}{1-\sin^2(x)}\]

OpenStudy (princessaurora):

oh okay

zepdrix (zepdrix):

Do you see how this might be useful for our problem?\[\large\rm \frac{\sin^2(x)}{\sqrt{\color{orangered}{1-\sin^2(x)}}}\]What can we do with the denominator here?

OpenStudy (princessaurora):

subtract cos^2(x)?

zepdrix (zepdrix):

subtract...? 0_o

OpenStudy (princessaurora):

add? sorry my brain is fried been working on this math stuff all day I'm homeschooled.

zepdrix (zepdrix):

We've shown, using our Pythagorean Identity, that 1-sin^2(x) is the same as cos^2(x). So we can simply `replace` our 1-sin^2(x) with cos^2(x), ya?

OpenStudy (princessaurora):

alright, I see

zepdrix (zepdrix):

That leaves us here,\[\large\rm \frac{\sin^2(x)}{\sqrt{\color{orangered}{\cos^2(x)}}}\]

zepdrix (zepdrix):

I guess we can proceed by taking the square root of that square,\[\large\rm =\frac{\sin^2(x)}{\cos(x)}\]And we'll have to do a little clever work at this point.

OpenStudy (princessaurora):

Is it tan^2(x)

OpenStudy (princessaurora):

simplified.

zepdrix (zepdrix):

Hmm, no. It's close to tan^2(x), but notice we don't have enough cosines in the bottom! :)

zepdrix (zepdrix):

Do you remember how tangent is defined in terms of sines and cosines? Bah I'll just type it out :)\[\large\rm \tan(x)=\frac{\sin(x)}{\cos(x)}\]

OpenStudy (princessaurora):

sin(x) tan(x)?

zepdrix (zepdrix):

So for our problem, maybe we can do something clever with the sines,\[\large\rm =\frac{\sin^2(x)}{\cos(x)}\quad=\frac{\sin(x)\cdot \sin(x)}{\cos(x)}\quad=\frac{\sin(x)}{\cos(x)}\cdot \sin(x)\]Replace the sin(x)/cos(x) by tan(x) to get sin(x)tan(x)? Yay good job \c:/

OpenStudy (princessaurora):

Yay thanks for putting up with my slow self..lol I appreciate the help a lot.

zepdrix (zepdrix):

np :3